OpenStudy (anonymous):

csc^2 θ = 8 cot θ + 10 solve the given equation

OpenStudy (anonymous):

can anyone help?

OpenStudy (anonymous):

Im here looking for answer for same problem :)

OpenStudy (anonymous):

divide by cot(theta)

OpenStudy (anonymous):

then simplify

OpenStudy (anonymous):

how do you divide by it?

OpenStudy (anonymous):

\[\csc^2/\cot=8 + 10/\cot ... \]

OpenStudy (anonymous):

\[\csc = 1/\cot so \csc^2/\cot = \csc^3 and 10/\cot = 10\csc\]

OpenStudy (anonymous):

i am confused? is that the answer to the equation? because it has to be somthing along the lines like 2pi/3 +2(pi)k

OpenStudy (anonymous):

no thats just how to do it. your gonna have to finish it.

OpenStudy (anonymous):

i dont know where to go from there?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

just plug stuff in its gotta be either 1/2 pi/2 ... exc

OpenStudy (loser66):

to me, I convert \[csc^2\theta = 1+cot ^2\theta\] therefore, yours is \[ 1+cot^2 \theta = 8cot \theta +10\\cot^2\theta-8cot\theta-9=0\\let ~\huge cot \theta =t \] \[t^2-8t -9 =0\] solve for t, then plug back to cot (theta) to find out the value of theta